Train cars are coupled together by being bumped into one another. Suppose two loaded train cars are moving toward one another. The first having a mass of 180,000 kg and a velocity of 0.300 m/s, and the second having a mass of 120,000 kg and a velocity of -0.120 m/s. What is their final velocity?
Answer #1

Answer -

Given,

Mass of train 1 = m1 = 180000 kg

Velocity of train 1 = V1 = 0.300 m/s

Mass of train 2 = m2 = 120000 kg

Velocity of train 2 = v2 = -0.120 m/s

Sign notation,

velocity from left to right is POSITIVE

and velocity from right to left is NEGATIVE.

So, train A is moving from left to right and train 2 is moving from right to left.

Now,

According to Conservation of momentum,

m1v1 + m2v2 = (m1 + m2)Vf

Vf = final velocity

Put the values,

(180000 kg * 0.300 m/s) + (120000 kg * -0.120 m/s) = (180000 kg + 120000) Vf

Vf = (180000 kg * 0.300 m/s) + (120000 kg * -0.120 m/s)/(180000 kg + 120000)

Vf = 0.132 m/s

Sign is positive So, motion is towards right.